OpenStudy (anonymous):

6/5+3/5x=21/10+3/2x+1/2 Clear all decimals or fractions first.

OpenStudy (anonymous):

multiply by 10 on both sides... 12 + 6x = 21 + 15x + 5 9x = -14 x = -14/9